OpenStudy (anonymous):

First on e answer and show work get a medal Divide 365 by 7 and write the anser as a mixed number. First on e answer and show work get a medal Divide 365 by 7 and write the anser as a mixed number. @Mathematics

OpenStudy (anonymous):

365 = 350+15=350+14+1=(7*50)+(7*2)+1=> 365/7 = 52 1/7